OpenStudy (gokuporter):

y=x^2-6x+5=0 Find the vertex and write as an ordered pair.

OpenStudy (gokuporter):

@jim_thompson5910

sammixboo (sammixboo):

You can find the vertex by finding the axis of symmetry, which can be found with this formula x = -b/2a This will help you find x. Once you have found your solution, plug in x into the original equation to find your y.

sammixboo (sammixboo):

Do you know how to find a and b?

OpenStudy (gokuporter):

No

OpenStudy (magentama202):

x=5, x=1 I believe this is the solution to the equation go to the link below for explanation... http://www.quickmath.com/webMathematica3/quickmath/equations/solve/basic.jsp#c=solve_stepssolveequation&v1=x%5E2-6x%2B5%3D0&v2=x just copy and paste into the url of a new tab and it should pop up sammixboo will probably be more help than me, so trust her/his answers...

sammixboo (sammixboo):

Well your original equation is set up in the form of \(\tt{y = ax^2 + bx + c}\) So now we know that a = 1 and b = -6!

OpenStudy (gokuporter):

@MAGENTAMA202 I want to learn how to find it, not just the answer.

OpenStudy (magentama202):

ok

sammixboo (sammixboo):

So we plug in our values for a and b into the axis of symmetry formula to find our x-coordinate for the vertex. \(\tt{x = \dfrac{-b}{2a}}\) \(\tt{x = \dfrac{6}{2(1)}}\)

OpenStudy (magentama202):

I'm just going to watch how sammixboo solved it if thats cool i havent spent much time on this type of questions before :)

sammixboo (sammixboo):

It's OK ;) If any of you guys have any further questions, let me know. Are you understanding everything so far, @Gokuporter

OpenStudy (gokuporter):

Yes

sammixboo (sammixboo):

OK, so do you know what \(\tt{x = \dfrac{6}{2(1)}}\) equals?

OpenStudy (gokuporter):

x=-3

sammixboo (sammixboo):

It's positive 3 :) Now we have found our axis of symmetry, we have also found the x-coordinate of our vertex. TO find our y-coordinate we must plug in 3 as x into the original equation, like so: \(\tt{y=3^2-6(3)+5}\)

sammixboo (sammixboo):

And solve. Do you know what this equals?

OpenStudy (gokuporter):

y=-4

OpenStudy (gokuporter):

@sammixboo

sammixboo (sammixboo):

Correct! So do you know what the vertex is?

OpenStudy (gokuporter):

(3,-4)

sammixboo (sammixboo):

Right :)